OpenStudy (anonymous):

Find S25 for 3 + 7 + 11 + 15 + …

OpenStudy (anonymous):

Solve the following system of equations: x − 2y = 14 x + 3y = 9

OpenStudy (anonymous):

what even zahabaha

OpenStudy (anonymous):

wat do u mean wat even??

OpenStudy (anonymous):

7-3=4 11-7=4 so it is AS which common difference = 4 Sum of AS =(first term + last term)( no. of terms ) /2 So, we have to find out the last term 7=3+4 11=3+2(4) 15=3+3(4) : : the 25th term = 3+24(4)=99 so sum = (3+99)(25)/2

OpenStudy (anonymous):

THANK YOU <3 you're perfect.